OpenStudy (anonymous):

Choose the correct simplification of (2xy7)2(y4)3.

OpenStudy (anonymous):

is 7 in power

OpenStudy (anonymous):

write this one correctly

OpenStudy (anonymous):

I dont get it?

OpenStudy (anonymous):

\[(2xy^7)^2(y^4)^3 or (2xy7)3(y4)3?\]

OpenStudy (anonymous):

\[(2xy^7)^{3(y^4)^3}?\]

OpenStudy (anonymous):

With the exponents

OpenStudy (anonymous):

The last one of the three?

OpenStudy (anonymous):

No the first one

OpenStudy (anonymous):

(2xy^7)^2(y^4)^3 = (4*x^2*y^14)(y^12) = 4x^2y^26

OpenStudy (anonymous):